# Keeping Up with the Models: Online Deployment and Routing of LLMs at Scale

### v1

@@ 5.2 Lower Bound
Consider an online LLM routing problem with $M\geq 2$ models over a horizon of $T$ queries. For any online routing algorithm $\mathcal{A}$, when $T$ is sufficiently large relative to $M$, its worst-case expected regret is bounded below by:
$\textnormal{\text{Regret}}(\mathcal{A})\;\geq\;\Omega(T^{2/3}).$
The $\Omega(T^{2/3})$ lower bound captures the intrinsic difficulty of continually tracking the best model as capabilities evolve. The proof uses a batch-based construction in which the “strong” model changes stochastically from one batch to the next, mirroring real LLM ecosystems where newer, better models apear regularly. An online algorithm must swiftly detect and exploit each transiently optimal model before it is displaced, and this necessity yields the unavoidable $T^{2/3}$ regret rate. Thus, the bound is not an artifact of the analysis but a fundamental limit for stage-based routing under dynamic model arrivals. The result also shows that system parameters—such as how often deployments occur and how many models can run concurrently—must be tuned carefully to approach this performance frontier.
Proof Sketch..
The proof constructs a class of adversarial problem instances designed to demonstrate that any online algorithm must incur significant regret on at least one of them. Specifically, the time horizon $T$ is partitioned into $N_{B}$ batches, each of equal length $\Delta=T/N_{B}$. In each batch $j\in\{1,\dots,N_{B}\}$, one out of $M$ models—denoted $m_{s_{j^{*}}}$—is designated as the strong model, offering an expected Bernoulli reward of $\mu_{strong}^{(j)}=\frac{1}{2}+j\epsilon$. The remaining $M-1$ models are treated as weak models, each yielding a lower expected reward of $\mu_{weak}^{(j)}=\frac{1}{2}+(j-1)\epsilon$. This construction captures a realistic abstraction of the modern LLM landscape, where new, more capable models continuously emerge while older ones remain accessible but less effective. The key statistical challenge lies in the unknown sequence of strong models, which must be discovered online. The reward gap $\epsilon>0$ plays a central role in the difficulty of the problem. To ensure that all mean rewards remain within the valid range $[0,1]$, the parameter $\epsilon$ is chosen such that $N_{B}\epsilon\leq c$ for constant $20<c<1/$ 2.
For a fixed batch $j$, the regret incurred in this batch is given by $R_{j}(P_{j,s_{j}^{*}})=\epsilon(\Delta-\mathbb{E}_{P_{j,s_{j}^{*}}}[N_{j,s_{j}^{*}}])$, where $N_{j,s_{j}^{*}}$ denotes the number of times the algorithm selects model $m_{s_{j}^{*}}$ during batch $j$. Leveraging information-theoretic arguments, which connect the difficulty of identifying the optimal model to the number of observations and the magnitude of the reward gap, we derive a bound on the total expected number of pulls across the set of strong models. This bound captures the algorithm’s limited ability to reliably discover and exploit the strong model within each batch, particularly when $\Delta$ is small and $\epsilon$ is narrow. We then lower bound the expected regret for batch $j$, under the assumption that the strong model index $s_{j}^{*}$ is chosen uniformly at random from $\{1,\dots,M\}$. This leads to the bound $R_{T}\gtrsim T\epsilon\left(\frac{1}{2}-\epsilon\sqrt{\Delta}\right)$. Optimizing the bound by selecting $\epsilon\asymp T^{-1/3}$ and $\Delta\asymp T^{2/3}$, and substituting these into the expression above, yields the desired lower bound. A detailed proof is provided in Appendix C. ∎
@@ 6 Experiments
This section presents the empirical evaluation of our proposed StageRoute algorithm, with a focus on its performance under diverse operational settings. All reported results are averaged over 10 independent simulation runs, each consisting of $T=10^{4}$ user queries. The experiments involve solving mixed-integer programming (MIP) subproblems using the academic version of the Gurobi Optimizer (version 12.0.1). All runs are executed on a machine equipped with a 12th Gen Intel(R) Core(TM) i9-12900HX processor.
Table 1: LLM comparison by release date: performance (Arena Score/Elo [13]) and output cost (per 1M tokens).
Model Performance Cost
GPT-4o 1336 \$10.00
Gemini-1.5-pro-002 1302 \$5.00
Yi-Lightning 1287 \$0.14
o1-mini 1304 \$0.60
Llama 3.3 70B Instruct 1257 \$0.40
DeepSeek-R1 1358 \$0.55
Gemini 2.0 Flash 1380 \$1.50
Claude 3.7 Sonnet 1300 \$15.00
Hunyuan-turbos 1296 \$0.28
Deepseek-v3 1373 \$0.28
Llama-4-Maveric 1417 \$0.82
GPT-4.1 1366 \$8.00
Grok-3-preview 1403 \$15.00
o3 by OpenAI 1413 \$40.00
Gemini-2.5-Pro 1446 \$10.00
6.1 Simulation Results
LLMs Pool. We consider a set of 15 LLMs, summarized in Table 1, ordered by their official release dates. For each model, we report both performance metrics  [13] and associated costs. Note that the naming of AI models can be highly nuanced, and multiple versions may exist under a similar label. For example, for Gemini 2.5 Pro, we used data corresponding to the Gemini-2.5-Pro-Preview-05-06 version. Additionally, performance scores may fluctuate over time due to model updates, shifts in the user voting population, or changes in evaluation benchmarks. Similarly, costs may vary across time or across versions. The data reported here corresponds to the specific versions and conditions used in our experiments.
These 15 models are introduced sequentially, with the first 6 available at time $t=1$ and one additional model becoming available every 1,000 rounds until all are accessible. Each model’s performance is normalized to lie within the $[0,1]$ interval. For each model $m$, we set its budget weight parameter $\alpha_{m}=0.4$, except for Yi-Lightning, for which we set $\alpha_{m}=1$. This is because Yi-Lightning is the cheapest model initially available to the algorithm, ensuring the feasibility of the mixed-integer optimization problem in every round. At each round, if the algorithm selects a model, it receives a noisy performance and cost observation, where the noise is sampled from a Gaussian distribution centered around the true value.
@@ 6.1 Simulation Results
Baselines. We compare our StageRoute against two baselines. The first is an oracle that has full knowledge of all relevant performance and cost statistics in advance and always selects the optimal set of models. This serves as an upper bound on achievable performance. The second baseline is a greedy strategy that, at each update point, selects the $M_{\max}$ models with the highest utility, computed as the ratio between the UCB of performance and the LCB of cost. This can in fact be viewed as a variant of a UCB algorithm, where selection is based on the UCB of the utility metric. For both our algorithm and the baselines, we set $\gamma=0.2$ and $b=2$. The cost constraint is enforced as a hard budget.
(a) Performance.
(b) Regret.
Figure 2: Performance with $M_{\max}=6$ and
Update Interval 250 rounds.
(c) 500 rounds.
(d) 1,000 rounds.
Figure 3: Varying update intervals with $M_{\max}$ fixed at 6.
@@
Figure 4: Varying $M_{\max}$ with update interval fixed at 500 rounds.
Figure 5: Model choice for $M_{\max}=3$ (Update interval at 500 rounds).
Overall Performance and Regret. We study the effect of two key hyperparameters on algorithm performance. The deployment update interval determines how frequently the algorithm can revise its set of actively deployed LLMs. We test intervals of 250, 500, and 1000 rounds. The second is maximum number of concurrently deployed models ($M_{\max}$), which limits the size of the active model pool at any time. We evaluate settings of $M_{\max}\in\{3,6,10\}$. Figures 55 illustrate the performance dynamics of the proposed algorithm under various settings. We begin with Figure 5, which shows the performance when $M_{\max}=6$ and the deployment update interval is 250 rounds. As shown in Figure 2(a), StageRoute attains performance that closely matches that of the optimal oracle. Meanwhile, Figure 2(b) shows that StageRoute incurs substantially lower regret than the greedy strategy, highlighting its superior ability to balance exploration and exploitation under uncertainty.
$\triangleright$ Effect of Deployment Update Interval. Figures 2(b), 2(c), and 2(d) illustrate the impact of different deployment update intervals. A shorter interval enables the algorithm to adapt more quickly to newly available LLMs as well as to evolving performance and cost estimates. In contrast, infrequent updates—such as every 1000 rounds, as shown in Figure 2(d)—slow down this adaptation process. This reduced responsiveness can lead to consistently higher regret across all algorithms.
$\triangleright$ Impact of $M_{\max}$. Figures 2(c), 3(a), and 3(b) show that a smaller value of $M_{\max}$ limits the algorithm’s flexibility in composing a diverse routing mix. In contrast, increasing the capacity to $M_{\max}=10$ (Figure 3(b)) allows both StageRoute and the greedy strategy to achieve strong performance. These results suggest that the advantage of StageRoute is most pronounced when deployment capacity is limited, while it remains competitive even in high-capacity regimes. Moreover, Figure 5 offers a concrete example of model selection behavior under $M_{\max}=3$ and an update interval of 500 rounds. The algorithm dynamically alternates between expensive proprietary models and cheaper open-source alternatives, effectively adapting to the evolving LLM landscape and maintaining performance in a cost-aware manner.
6.2 Real-World Dataset Evaluation
We then present experimental results on a real-world dataset. Our evaluation is based on the Router-Bench dataset [24], a comprehensive benchmark comprising 36,497 queries sampled from eight diverse natural language processing (NLP) datasets. These datasets encompass both Chinese and English languages and cover a broad spectrum of core tasks relevant to LLMs. Specifically, the benchmark includes seven established evaluation datasets, along with an additional dataset targeting retrieval-augmented generation (RAG).
The seven core datasets span the following categories: commonsense reasoning, including HellaSwag [46], WinoGrande [37], and ARC Challenge [15]; knowledge-based language understanding, represented by MMLU [23]; open-domain dialogue, using MT-Bench [48]; mathematical reasoning, evaluated through GSM8K [16]; and code generation, based on MBPP [8]. The additional RAG dataset is designed to reflect realistic retrieval-augmented tasks encountered in production environments. For each query in the dataset, responses from 11 distinct LLMs are provided, together with corresponding performance scores and cost measurements.
Model Avg Performance Avg Cost
claude-instant-v1 0.5900 \$0.001236
claude-v1 0.6480 \$0.005870
claude-v2 0.5116 \$0.006153
meta/llama-2-70b-chat 0.6059 \$0.001337
WizardLM/WizardLM-13B-V1.2 0.5392 \$0.000142
meta/code-llama-instruct-34b-chat 0.5040 \$0.000550
mistralai/mistral-7b-chat 0.4999 \$0.000139
gpt-3.5-turbo-1106 0.6867 \$0.000709
gpt-4-1106-preview 0.8048 \$0.007943
zero-one-ai/Yi-34B-Chat 0.7153 \$0.000558
mistralai/mixtral-8x7b-chat 0.6504 \$0.000414
Figure 6: Cumulative performance score over time on the real-world dataset.
In our experiments, we simulate a total of $T=36,\!497$ rounds. At each round, a query is sampled uniformly at random from the dataset, including its associated prompt. The algorithm then selects a model to serve the query, and subsequently receives a noisy reward signal (or correctness label) along with the corresponding cost.
Initially, only 5 models are available for deployment. Then, for every additional 5,000 queries, one new model becomes eligible for deployment, following the release-date ordering specified above. For the remaining parameters, we set $\gamma=0.2$ and $b=0.0002$. The cumulative reward achieved by different methods under this setting is illustrated in Figure 6. As shown in the figure, our algorithm consistently outperforms the Greedy strategy by an even larger margin compared to the synthetic dataset experiments. This further demonstrates the effectiveness and robustness of our method in real-world scenarios.
Limitations and Future Work
In this paper, we introduced a stage-based framework for online LLM routing that explicitly handles changing model availability under budget, capacity, and deployment constraints. The resulting algorithm balances exploration and exploitation in this challenging setting, and we supported it with provable regret bounds and strong empirical results. Regarding societal impacts, our framework has the potential to facilitate broader adoption of LLMs by lowering deployment costs and improving energy efficiency. However, increased accessibility and proliferation of LLMs may also give rise to negative societal consequences. Several limitations remain. Our reward and cost estimates rely on UCB/LCB statistics derived from past observations, which may be inadequate in scenarios where a model’s performance or pricing varies significantly with the input context. In fact, if the number of deployable models is smaller than the number of distinct context clusters, any routing policy suffers $\Omega(T)$ worst-case regret. Future work can exploit structure in the context space. For instance, if rewards or costs vary linearly with contextual features, more sophisticated contextual bandit techniques may yield tighter regret guarantees and better practical performance. Exploring such structured settings is an appealing direction for further research.
Appendix A Summary of Notation

### v2

@@ 4.2 Lower Bound
For any online policy $\mathcal{A}$ and any choice of update frequency $K$ and concurrency cap $M_{\max}$, there exists a stochastic, piecewise-stationary LLM-routing instance such that the expected regret against the time-varying oracle satisfies
$\mathrm{Regret}(\mathcal{A})\geq\Omega(T^{2/3}).$
This $\Omega(T^{2/3})$ bound captures the intrinsic difficulty of continually tracking the best model as capabilities evolve. The construction mirrors real LLM ecosystems: in each batch a (newer) model is marginally stronger than the rest, and the identity of the strongest model changes over batches. Importantly, the baseline level of rewards also drifts upward over time, reflecting that even “weaker” new releases can outperform old ones. This is unlike classic lower bounds that keep suboptimal arms at a fixed mean (e.g., $1/2$) and change only the identity of the best arm.
Proof ideas and Intuition. We construct a family of “streaming” instances that mimics a live LLM marketplace: the time horizon is split into approximately $T^{1/3}$ epochs, in each batch a different model is slightly better than the others, and the whole performance frontier drifts upward across batches (so even “weaker” newcomers can surpass yesterday’s best). We choose the batch length and performance gaps so that any algorithm cannot reliably identify it with the limited information available before the epoch ends. Because the identity of the best model changes next batch, information gained earlier quickly goes stale. Any policy is thus forced into repeated “partial discovery”, incurring a nontrivial loss in each batch, and summing over all batches yields total regret on the order of $T^{2/3}$. Full details are given in Appendix D.
New technical elements. Two aspects differ from standard MAB/CMAB lower bounds: (i) we do not use a fixed baseline where only the identity of the best arm flips. Here the entire frontier drifts, matching LLM practice; and (ii) the hardness persists even if the system can redeploy every round and keep all models live, so the rate is intrinsic to tracking an evolving frontier, not an artifact of staging or capacity limits.
Discussion. Theorem 2 conveys a practical message: even with aggressive adaptivity and large live sets, there is a fundamental rate limit on how quickly a system can keep up as the performance frontier shifts. Our upper bound matches this lower bound up to logarithmic factors with respect to the number of queries $T$, establishing near-optimality.
(a) The decision heatmap of StageRoute.
(b) The decision heatmap of the optimal Oracle.
Figure 2: Comparison of decision heatmaps for StageRoute and the Oracle with $M_{\max}=5,b=0.001$, update interval=1000. Darker colors indicate higher selection probabilities.
@@ 5 Experiments
Datasets and Candidate LLMs. We evaluate on RouterBench (Hu et al., 2024), covering 36,497 queries across eight datasets in English and Chinese (commonsense, knowledge, dialogue, math, code, and RAG). Each query includes responses from 11 LLMs with per-query scores and costs. Full dataset descriptions and the model list appear in Appendix E.
@@
Baselines. We compare StageRoute against three baselines. The first is an oracle that, with full knowledge of all performance and cost statistics, always selects the optimal deployment set, serving as an upper bound on achievable performance. The second is a greedy strategy that, at each update point, selects the $M_{\max}$ models with the highest utility, computed as the ratio between the UCB of performance and the LCB of cost. This approach can be viewed as a variant of a UCB algorithm where selection is based on the UCB of the utility metric. The third baseline is a uniform sampling strategy, which randomly selects models for deployment and may substantially exceed the budget. To our knowledge, no existing methods are specifically designed for this LLM deployment problem.
Figure 3: Cumulative regret.
Implementation Details. We simulate a total of T=36,497T=36,\!497 rounds. In each round, a query is sampled uniformly at random from the dataset. The algorithm then selects a model to serve the query and subsequently receives the performance score and associated cost. Initially, 5 models are available. Thereafter, for every 5,000 queries, a new model becomes eligible for deployment, following the release-date ordering in Table 3. We set the confidence parameter $\gamma=0.1$. All reported results are averaged over 10 independent runs. The experiments involve solving mixed-integer programming (MIP) subproblems using the Gurobi Optimizer (v12.0.1, academic license) on a machine equipped with a 12th Gen Intel(R) Core(TM) i9-12900HX processor.
Computational Overhead. Our two-stage design keeps the computational overhead low, allowing the entire experimental run to complete in under 10 minutes. The deployment MIP is solved only infrequently on small instances (taking sub-seconds), while the per-query routing involves a tiny LP over the active set that executes in milliseconds. Therefore, using either Gurobi or an open-source solver (e.g., CBC, GLPK, HiGHS) is feasible. Furthermore, since parameters change only slightly between iterations, we can leverage warm-starting to further accelerate computation. For the routing LP, we can reuse the previous basis or solution; similarly, for the deployment MIP, we can provide the prior active set as an initial feasible solution (MIP start), given the minimal changes to the candidate pool (Zhang et al., 2025a).
Overall Performance. We first present the main results under a representative setting ($M_{\max}=5,b=0.001$, update interval=1000), and then conduct a detailed sensitivity analysis across key hyperparameters. Figure 3 shows the cumulative regret under this default configuration. Across all settings we test (see Figure 5 for a full overview), our algorithm exhibits consistently slow regret growth, substantially outperforming the baselines. Notably, while the uniform sampling strategy appears to outperform the greedy baseline in some cases, this is an artifact of its tendency to significantly overspend the budget, an issue we will analyze further in the performance-cost evolution.
Optimal Model Set Identification. As our work emphasizes the importance of deployment, we first analyze whether StageRoute can identify the optimal model set. Figure 2 compares the decision probabilities of StageRoute and the Oracle under the representative setting mentioned above. The horizontal axis represents deployment intervals, while the vertical axis (bottom to top) corresponds to the model arrival order. It is evident that when a new model arrives, StageRoute initially explores it before quickly converging to the new optimal model set, closely mirroring the Oracle’s behavior. This confirms that our deployment strategy is effective at tracking the optimal available model set.
Figure 4: Performance-cost evolution of different algorithms.
Performance-Cost Evolution. To further validate our algorithm’s efficiency, Figure 4 illustrates the performance-cost trajectory for each algorithm, again under the same representative setting. Colors transition from blue (initial stages) to red (final stages). The figure shows that, except during initial exploration and periods when new models arrive, the operating points of our algorithm closely track those of the Oracle. In contrast, the greedy strategy proves overly conservative, while uniform sampling consistently violates the budget for suboptimal performance. These observations reinforce our central claim: selecting a high-quality set of models for deployment is fundamental to achieving efficient routing, and StageRoute successfully balances high performance with strict budget adherence.
Sensitivity Analysis. We now analyze StageRoute’s sensitivity to key hyperparameters.
(1) Impact of $M_{\max}$. Figures 3, 5(a), and 5(b) present results for different $M_{\max}$ values under a budget of $b=0.001$ and an update interval of 1000. The results demonstrate that StageRoute adapts well to this parameter, maintaining robust performance across all settings.
(2) Effect of Deployment Update Interval. Figures 3, 5(c), and 5(d) illustrate the impact of varying the deployment update interval with $M_{\max}=5$ and $b=0.001$. An interval of 1000 rounds yields the lowest regret, highlighting the importance of selecting an appropriate update frequency.
(3) Effect of Budget Constraint. Figures 3 and 5(e) compare performance under different budget constraints. Counterintuitively, a more relaxed budget leads to higher regret. This phenomenon can be attributed to two factors. First, a larger budget also raises the performance of the Oracle, making the benchmark more challenging. Second, we use a fixed confidence radius $\gamma$ for all settings; in practice, increasing $\gamma$ in proportion to the budget may be beneficial.
Extending to State-of-the-Art Models. To verify that our StageRoute framework applies to the latest, most powerful models, we conduct additional simulations incorporating recent LLMs. These results, detailed in Appendix E, confirm that StageRoute continues to achieve minimal regret.
@@
(c) 500 rounds.
(d) 2000 rounds.
(e) b=0.002.
Figure 5: Cumulative regret under varying hyperparameters. The default setting is $M_{\max}=5$, update interval = 1000 rounds, and $b=0.001$ (Figure 3).
6 Conclusion
In this paper, we introduced StageRoute, a novel framework for online LLM deployment and routing. We are the first to formalize this problem in a dynamic setting with streaming LLM model arrivals, addressing the challenge of selecting an optimal deployment set under a strict concurrency cap. Our algorithm manages deployment at discrete stages while tactically routing queries in real time, respecting both budget and per-model throughput limits. We established the near-optimality of our algorithm with theoretical analysis, including matching upper and lower bounds, and demonstrated its practical effectiveness through extensive experiments on real-world benchmarks.
Acknowledgements
This work was supported in part by the National Science Foundation (NSF) grants 2148309, 2315614 and 2337914, and the National Institutes of Health (NIH) grant 1R01HL184139-01, and was supported in part by funds from OUSD R&E, NIST, and industry partners as specified in the Resilient & Intelligent NextG Systems (RINGS) program. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.
Ethics Statement
This research focuses on the operational efficiency of LLM systems. By making deployment and routing more cost-effective, our work can broaden access to AI technologies and reduce energy consumption. However, we acknowledge that increased accessibility may also lower the barrier for malicious use of LLMs. Our framework does not mitigate the inherent risks of language models, such as bias or misinformation generation, and should be implemented alongside robust safety and content moderation protocols.
Appendix A Related Work
A.1 LLM Routing
The central aim of LLM routing is to strike the best balance between task performance (e.g., response quality or accuracy) and operational metrics such as cost and latency (Ding et al., 2024; Aggarwal et al., 2024). Existing work follows three main architectural patterns. Ensemble strategies query multiple models in parallel to boost robustness, at the expense of higher cost and latency Wang et al. (2023a); Jiang et al. (2023). Cascade strategies issue queries sequentially—typically starting with a cheaper model and escalating only when necessary—thereby reducing cost but sometimes increasing latency (Chen et al., 2023; Gupta et al., 2024; Yue et al., 2024; Aggarwal et al., 2024).Direct‐routing strategies train a policy or classifier that selects a single LLM per query (Ong et al., 2025; Feng et al., 2025; Zhang et al., 2025b; Zhuang et al., 2025). Benchmark suites such as RouterBench (Hu et al., 2024) facilitate systematic comparison of these approaches. Related work on mixture-of-experts (MoE) models explores routing within a single large model (Du et al., 2022; Fedus et al., 2022; Riquelme et al., 2021). More recently, bandit formulations have been applied to static LLM routing (Wang et al., 2025; Dai et al., 2024; Li, 2025; Nguyen et al., 2024; Poon et al., 2025). Most prior studies, however, assume a fixed set of available models and focus solely on per-query decisions. In contrast, our work model a dynamic model pool with streaming arrivals and introduce staged deployment updates, where the active model set is subject to online selection under a strict concurrency cap, cost budget, and throughput limits. To our knowledge, we are the first to formalize and solve this more realistic and challenging problem.
A.2 Multi-armed Bandits
Our formulation builds on the multi-armed-bandit (MAB) paradigm, where an agent maximizes its cumulative payoff through exploration and exploitation in online environment (Auer et al., 2002; Slivkins, 2019). Three MAB extensions are especially pertinent: (i) Bandits with knapsacks (BwK). Here each arm pull yields a reward and consumes limited resources from one or more budgets; the objective is to maximize total reward without overspending (Badanidiyuru et al., 2013; Agrawal and Devanur, 2014; Immorlica et al., 2019; Kesselheim and Singla, 2020; Bernasconi et al., 2024; Guo and Liu, 2025). Our long-term cost constraint fits naturally into this framework. (2) Streaming bandits. In this setting new arms arrive over time—often under memory or attention limits—so the agent must adapt to a continually expanding action set (Assadi and Wang, 2020; Jin et al., 2021; Agarwal et al., 2022; Wang, 2023; Li et al., 2023; Shao and Fang, 2025; Zhu and Huang, 2025). The steady appearance of new LLMs places our problem squarely in this category. (3) Combinatorial Multi-Armed Bandits (CMAB). The algorithm faces a fixed, known set of base arms from which superarms (subsets) are chosen in each round (Cesa-Bianchi and Lugosi, 2009; Chen et al., 2013; Qin et al., 2014; Li et al., 2016; Chen et al., 2018; Liu et al., 2024; 2025).
The core distinctions arise from the two-level structure and unique constraints inherent to the realistic online LLM deployment and routing problem. Standard models like BwK and streaming bandits lack the combinatorial selection. While CMAB addresses superarm selection, it is fundamentally misaligned with our problem’s dynamics: it assumes a static set of base arms and makes per-round decisions, whereas our core challenges are a dynamic model pool and staged, irreversible commitment, where a deployed set remains fixed for a long duration. The regret is thus a function not only of the chosen set but also of the tactical routing policy executed over thousands of queries within that stage. This stateful, hierarchical structure is beyond the scope of traditional bandit formulations. Due to these fundamental differences, existing algorithms and regret analyses are not applicable. Our work bridges this gap by developing a new framework and novel analytical tools tailored to the unique challenges of online LLM deployment and routing problem.
Appendix B Summary of Notation